# pre-trained model AST

### v1

@@ 6.5. RQ4: Comparison to SOTA
Weights-Only Unpickler (PyTorch, 2024c) 84 157 95 0 37.6% 0%
@@
We compare PickleBall with three existing state-of-the-art (SOTA) tools that share the same goal of defending against our threat model described in §4. As discussed in §2.2, existing pickle model defense tools fall into two categories: model scanners and restricted loading environments (like PickleBall). We compare against two model scanners: ModelScan (AI, 2024a), and the scanner implemented by Casey et al. (Casey et al., 2024) (henceforth, ModelTracer). [^7] The authors provided access to the tool for evaluation purposes. ModelScan is a static analysis tool that applies a rigid denylist to make determinations about models, and is integrated into Hugging Face. ModelTracer is a dynamic analysis tool that traces the model’s invocations while it is loaded via pickle.loads() and torch.load(), and similarly applies a rigid denylist to make its determinations.
@@ 6.5.1. Methods
We evaluate the model scanning tools by providing each model in our dataset as an input to the tool. We expect the model scanners to alert when provided a malicious model input, and otherwise not to alert. We consider correctly identified malicious models as true positives, correctly identified benign models as true negatives, incorrectly identified malicious models as false negatives, and incorrectly identified benign models as false positives.
We evaluate the weights-only unpickler by attempting to load each model in our dataset using the PyTorch loading API with the weights-only unpickler enabled. We use the weights-only unpickler’s default policy while loading models. We expect restricted loading environments like the weights-only unpickler and PickleBall to correctly load benign models and to raise exceptions when loading malicious models. For parity when comparing with the model scanning tools, we consider raising an exception during malicious model loading as a true positive, correctly loading a benign model as a true negative, incorrectly rejecting a benign model as a false positive, and incorrectly loading a malicious model as a false negative. We apply the same criteria to compare PickleBall with these tools, using the dataset and results obtained in Section˜6.2 and Section˜6.3.
@@ 6.5.2. Results
RQ4 Summary: While PickleBall prevents all malicious models from loading, model scanning tools fail to identify all malicious models. The weights-only unpickler is also effective at preventing malicious models from loading, but is less effective than PickleBall at loading benign models due to its rigid default policy.
@@ 7. Discussion and Future Work
PickleBall’s remaining attack surface: PickleBall reduces the available attack surface by significantly restricting access to callables, but it does not guarantee that the remaining callables cannot be composed in a malicious payload (§5.3). Related software security approaches provide desirable-yet-imperfect reductions in attack surfaces, whether by removing access to system calls (DeMarinis et al., 2020; Gaidis et al., 2023) or executable gadgets and control flows (van der Veen et al., 2015; Zhang et al., 2013; Agadakos et al., 2019). Similarly, PickleBall could be compromised by code reuse or property-oriented programming-like techniques (Park et al., 2022; Dahse et al., 2014) to stitch together permitted calls to construct an exploit. Thus far, no such attacks have been observed, but the question remains: can we generate malicious payloads that obey the policy constraints enforced by PickleBall and the weights-only unpickler?
Huang et al. (Huang et al., 2022) studied manual implementations of restricted unpicklers in the general (non-ML) setting, and devised attack strategies to try to overcome callable allowlists. When faced with well-implemented restricted unpicklers (i.e., when recursive attribute look-ups and indexing are disallowed by design, as in PickleBall and the weights-only unpickler), their approach degrades to manual policy inspection. Future work should identify properties of callables to automatically distinguish whether a callable can be used maliciously.
Long-term outlook for pickle in ML: Pickle remains a popular model format (§3), despite more secure alternatives. Each major model format provides tradeoffs in flexibility (pickle), security (SafeTensors), and efficiency (GGUF). Pickle is flexible, as it can serialize virtually any Python object, including complex models with custom operators and non-standard data structures. SafeTensors was developed for security-sensitive deployments, with a structured, memory-mapped format. GGUF maximizes performance on inference-optimized runtimes. Because these formats are complementary and used as defaults in different popular frameworks and ecosystems, we expect them to coexist going forward. PickleBall does not discourage the adoption of secure alternatives to pickle, but provides a secure option for the (as our data show, large and growing) pickle population.
@@
Generalizing PickleBall: PickleBall is designed to protect pickle model loading and is evaluated on models found on Hugging Face, but its approach can generalize to protect (1) other model formats; and (2) other pickle applications.
PickleBall aims to protect pickle deserialization in the context of ML models, but its policy generation and enforcement modules do not rely on any ML-specific properties. This approach allows PickleBall to protect other applications that receive pickle data. To generalize, PickleBall’s analysis requires that the intended type of the pickle object is known before loading (§5.1). In the ML setting, this is reasonable because the protected program is a client application. Other approaches are needed when the security analysis does not know a priori the intended type of the serialized object (David et al., 2024; Zhang et al., 2024).
PickleBall’s approach works for the pickle format because it has (dangerously) expressive deserialization operations, and is used by trusted libraries that implement their own custom model classes — which are analyzed to learn a specification for allowed behaviors. Other model formats that meet these criteria are candidates for protection in the PickleBall approach. Zhu et al. show that the TensorFlow SavedModel format has undesirable operations (Zhu et al., 2025); libraries that extend the TensorFlow model class with their own custom behaviors could use a PickleBall approach to restrict the allowed behaviors, but we are not aware of any that do. Formats like SafeTensors and GGUF are not known to have dangerous operations; if any were discovered, then a PickleBall approach might apply for identifying when to permit certain operations. We aim to explore which other model formats meet these criteria for PickleBall to assist in securely loading.
We evaluated PickleBall using models sourced from Hugging Face, but PickleBall will work similarly for models from any platform. The inputs to PickleBall are ML libraries and pickle models, which are ML artifacts that are not tied to the hosting platform. Hugging Face is the largest model hosting platform, with over 1.8M models available in July 2025. ModelScope (mod, 2025) is a recent hub managed by China’s Alibaba and hosts 80K models. It imitates Hugging Face’s design and likewise has models with varying serialization formats. Other model communities, including Qualcomm AI Hub (qua, 2025), PyTorch Hub (pyt, 2025), and TensorFlow Hub (ten, 2025), have fewer than 500 models each and many are also hosted on Hugging Face. The ONNX Model Zoo (onn, 2025) is now deprecated and archived on Hugging Face. Hugging Face models are representative of the kinds of models that PickleBall can harden.
@@
PickleBall generates policies quickly, completing in under 30 seconds for each library in our evaluation (see §6.4, Figure˜6). This is reasonable for a task that must only occur when the library changes, and not every time PickleBall loads a model.
@@
PickleBall provides opportunities for easier policy maintenance when it is adopted further upstream in the model development life cycle. PickleBall enables model consumers to protect themselves from malicious model producers by trusting only code from library maintainers. Library maintainers can adopt PickleBall themselves to generate policies automatically, as part of the library release process, and provide the updated policies alongside the libraries.
@@ 8. Related Work
ML Model Loading Security: Pickle is not intended for untrusted data, but its proliferation as a model format created a security problem. To bring attention to the issue, security company Trail of Bits released the fickling tool for manipulating and analyzing pickle programs in ML models (of Bits, 2024). New proposals for identifying malicious pickle models include dynamic (Casey et al., 2024) and static (Zhao et al., 2024) scanners. Scanners take a model as input, and attempt to make an assessment of it based on fixed rules about malicious behaviors. Instead, PickleBall takes a model and the source code library that allegedly produced the model for context, and produces policies for the model based on that context. We showed in Section˜6.5 that PickleBall’s tailor-made policies result in no false negatives while comparative model scanners do produce false negatives due to their fixed rules.
Deserialization Attacks and Defenses: Deserialization vulnerabilities exist beyond the ML context. PainPickle (Huang et al., 2022) explored Python pickle security by creating a taxonomy of errors in custom Unpickler implementations, and devised attack strategies. We use their contributions to guide the proper implementation of PickleBall’s loader.
Other programming languages have also have insecure deserialization APIs that need to be secured. Quack (David et al., 2024) propose a generic deserialization defenses for PHP by employing a “static duck typing” static analysis, and Zhang et al. (Zhang et al., 2013) propose a static analysis defense for Java. Python’s pickle deserialization is a more expressive interface than in PHP or Java, which are unable to directly invoke functions; this expressivity adds complexity to security policies.
@@ 9. Conclusion
Serialization and deserialization enable code and data exchange, enabling software re-use. Many recent works observe security vulnerabilities in deserialization, across various programming languages and contexts. We specifically examined the security vulnerabilities in pre-trained model deserialization that result from the use of (dynamically typed) Python and the reliance on Python’s (insecure) pickle format. We found that pickle use is common among the most popular models on Hugging Face, and that existing defenses are insecure or inapplicable to a substantial fraction of these models. Our PickleBall approach applies a novel program analysis to add greater type-safety to model deserialization. In our evaluation, PickleBall supported most existing benign models transparently while preventing all known attacks in malicious models. We believe PickleBall is a promising complement to existing security resources in the pre-trained model ecosystem.
Data Accessibility
Upon acceptance, we will open-source PickleBall, our evaluation scripts, and our pickle +ML security dataset.
@@
Our primary contribution is PickleBall, a defense that improves security for ML models. Our Hugging Face study abides by the platform policies for API use. We did identify some possible shortcomings of existing security approaches, and have followed a responsible disclosure process—we privately disclosed the ModelScan bypass described in Appendix B to its author, the company Protect AI, and similarly to the ModelTracer team.
@@ Acknowledgements
We thank our anonymous reviewers for their valuable feedback. We thank Joseph Lucas, John Irwin, and Suha Hussain for their insightful discussions of pickle and model security. We thank Beatrice Casey for providing access to the ModelTracer tool (Casey et al., 2024), and to Shenao Wang for sharing malicious models discovered on Hugging Face (Zhao et al., 2024). Andreas Kellas is supported by the Department of Defense (DoD) National Defense Science and Engineering (NDSEG) Fellowship program. James Davis is supported in part by funding from Socket. Junfeng Yang is supported in part by funding from Google.

### v2

@@ 6.5. RQ4: Comparison to SOTA
Weights-Only Unpickler (PyTorch, 2024c) 84 157 95 0 37.7% 0%
@@
We compare PickleBall with three existing state-of-the-art (SOTA) tools that share the same goal of defending against our threat model described in Section 4. As discussed in Section 2.2, existing pickle model defense tools fall into two categories: model scanners and restricted loading environments (like PickleBall).
We compare against two model scanners: ModelScan (AI, 2024a), and the scanner implemented by Casey et al. (Casey et al., 2024) (henceforth ModelTracer). [^7] The authors provided access to the tool for evaluation purposes. ModelScan is a static analysis tool that applies a denylist to make determinations about models, and is integrated into Hugging Face. ModelTracer is a dynamic analysis tool that traces the model’s invocations while it is loaded via pickle.loads()/torch.load(), and similarly applies a denylist.
@@ 6.5.1. Methods
We evaluate the model scanning tools by providing each model in our dataset as an input to the tool. We expect the model scanners to alert when provided a malicious model input, and otherwise not to alert.
We consider correctly identified malicious models as true positives, correctly identified benign models as true negatives, incorrectly identified malicious models as false negatives, and incorrectly identified benign models as false positives.
We evaluate the weights-only unpickler by attempting to load each model in our dataset using the PyTorch loading API with the weights-only unpickler enabled. We use the weights-only unpickler’s default policy while loading models. We expect restricted loading environments like the weights-only unpickler and PickleBall to succeed when loading benign models and to raise exceptions when loading malicious models.
For parity when comparing with the model scanning tools, we consider raising an exception during malicious model loading as a true positive, correctly loading a benign model as a true negative, incorrectly rejecting a benign model as a false positive, and incorrectly loading a malicious model as a false negative.
@@ 6.5.2. Results
RQ4 Summary: While PickleBall prevents all malicious models from loading, model scanning tools fail to identify all malicious models. The weights-only unpickler is also effective at preventing malicious models from loading, but is less effective than PickleBall at loading benign models due to its limited default policy.
@@ 7. Discussion and Future Work
PickleBall’s remaining attack surface: PickleBall reduces the available attack surface by significantly restricting access to callables, but it does not guarantee that the remaining callables cannot be composed in a malicious payload (§5.3). PickleBall could be compromised by code reuse techniques (Park et al., 2022; Dahse et al., 2014) to stitch together permitted calls to construct an exploit. Thus far, no such attacks have been observed, but the question remains: can we generate malicious payloads that obey the policy constraints enforced by PickleBall and the weights-only unpickler?
Huang et al. (Huang et al., 2022) studied manual implementations of restricted unpicklers in the general (non-ML) setting, and devised attack strategies to overcome callable allowlists. When faced with well-implemented restricted unpicklers (i.e., when recursive attribute look-ups and indexing are disallowed by design, as in PickleBall and the weights-only unpickler), their approach degrades to manual policy inspection. Liu et al. (Liu et al., 2025) implemented PickleCloak to automatically detect useful pickle gadgets, but apply it to bypass scanner deny-lists rather than restricted loader allow-lists. Future work should identify properties of callables to automatically distinguish whether a callable can be used maliciously to bypass restricted allow-lists.
Long-term outlook for pickle in ML: Pickle remains a popular model format (§3), despite more secure alternatives. Each major model format provides tradeoffs in flexibility (pickle), security (SafeTensors), and efficiency (GGUF). Pickle is flexible, as it can serialize virtually any Python object, including complex models and non-standard data structures. SafeTensors was developed for security-sensitive deployments, with a structured, memory-mapped format. GGUF maximizes performance on inference-optimized runtimes. Because these formats are complementary and used as defaults in different popular frameworks and ecosystems, we expect them to coexist going forward. PickleBall does not discourage the adoption of secure alternatives to pickle, but provides a secure option for the large and growing pickle population.
@@
Generalizing PickleBall: PickleBall is designed to protect pickle model loading and is evaluated on models found on Hugging Face, but its approach can generalize to protect (1) other model formats; and (2) other pickle applications. PickleBall aims to protect pickle deserialization for ML models, but its approach does not rely on any ML-specific properties.
This approach allows PickleBall to protect other applications that receive pickle data. To generalize, PickleBall’s analysis requires that the intended type of the pickle object is known before loading (§5.1). In the ML setting, this is reasonable because the protected program is a client application. Other approaches are needed when the security analysis does not know a priori the intended type of the serialized object (David et al., 2024; Zhang et al., 2024).
PickleBall’s approach works for the pickle format because it has (dangerously) expressive deserialization operations, and is used by trusted libraries that implement their own custom model classes. Other model formats that meet these criteria are candidates for protection in the PickleBall approach. Zhu et al. show that the TensorFlow SavedModel format has undesirable operations (Zhu et al., 2025); libraries that extend the TensorFlow model class with their own custom behaviors could use a PickleBall approach to restrict the allowed behaviors, but we are not aware of any that do. Formats like SafeTensors and GGUF are not known to have dangerous operations; if any were discovered, then a PickleBall approach might apply for identifying when to permit certain operations. We aim to explore which other model formats meet these criteria for PickleBall to assist in securely loading.
We evaluated PickleBall using models sourced from Hugging Face, but PickleBall will work similarly for models from any platform. The inputs to PickleBall are ML libraries and pickle models, which are ML artifacts that are not tied to the hosting platform. Hugging Face is the largest model hosting platform, with over 1.8M models available in July 2025. ModelScope (mod, 2025) is a recent hub managed by China’s Alibaba and hosts 80K models. It imitates Hugging Face’s design and likewise has models with varying serialization formats. Other model communities, including Qualcomm AI Hub (qua, 2025), PyTorch Hub (pyt, 2025), and TensorFlow Hub (ten, 2025), have fewer than 500 models each and many are also hosted on Hugging Face. The ONNX Model Zoo (onn, 2025) is now deprecated and archived on Hugging Face. Hugging Face models are representative of the kinds of models that PickleBall defends against.
@@
PickleBall generates policies quickly, completing in under 30 seconds for each library in our evaluation (§6.4, Figure˜6). This is reasonable for a task that must only occur when the library changes, and not every time PickleBall loads a model.
@@
PickleBall provides opportunities for easier policy maintenance if it gets adopted further in the model development life cycle. When library maintainers adopt PickleBall to generate policies automatically in the library release process, they can provide the updated policies alongside the libraries directly to users.
@@ 8. Related Work
ML Model Loading Security: Pickle is not intended for untrusted data, but its proliferation as a model format created a security problem. To bring attention to the issue, security company Trail of Bits released the fickling tool for manipulating and analyzing pickle programs in ML models (of Bits, 2024). The fickling module provides manually-crafted allow-lists for select ML libraries, similarly to PickleBall, but requires expert-led audits to maintain the allow-lists, where PickleBall aims for automatic policy creating. New proposals for identifying malicious pickle models include dynamic (Casey et al., 2024) and static (Zhao et al., 2024) scanners. Scanners take a model as input, and attempt to make an assessment of it based on fixed rules about malicious behaviors. Instead, PickleBall takes a model and the source code library that allegedly produced the model for context, and produces policies for the model based on that context. We showed in Section˜6.5 that PickleBall’s tailor-made policies result in no false negatives while comparative model scanners do produce false negatives due to their fixed rules.
Deserialization Attacks and Defenses: Deserialization vulnerabilities exist beyond the ML context. PainPickle (Huang et al., 2022) explored Python pickle security by creating a taxonomy of errors in custom Unpickler implementations, and devised attack strategies. We use their contributions (and suggestions) to guide the proper implementation of PickleBall’s loader.
Other programming languages also have insecure deserialization APIs that need to be secured. Quack (David et al., 2024) proposes a generic deserialization defenses for PHP by employing a “static duck typing” static analysis, and Zhang et al. (Zhang et al., 2024) propose a static analysis defense for Java. Python’s pickle deserialization API is more expressive than those in PHP or Java, which are unable to directly invoke functions; this expressivity adds complexity to security policies.
@@ 9. Conclusion
Serialization and deserialization enable code and data exchange. Many recent works observe security vulnerabilities in deserialization, across various programming languages and contexts. We specifically examined the security vulnerabilities in pre-trained model deserialization that result from the use of (dynamically typed) Python and the reliance on Python’s (insecure) pickle format. We found that pickle is common among the most popular models on Hugging Face, and that existing defenses are insecure or inapplicable to a substantial fraction of these models. Our PickleBall approach applies a novel program analysis to add greater security to model deserialization. In our evaluation, we demonstrated that PickleBall supported most existing benign models while preventing all known attacks in malicious models. We believe PickleBall is a promising complement to existing security resources in the pre-trained model ecosystem.
@@
We considered the stakeholders involved in this study (Davis et al., 2025) and believe our results offer a net benefit. Our primary contribution is PickleBall, a defense that improves security for ML models. Our Hugging Face study abides by the platform policies for API use. We identified some shortcomings of existing defenses, and we followed a responsible disclosure process to report them.
Data Availability
The PickleBall source code is maintained at: https://github.com/columbia/pickleball. Our complete evaluation artifact is made available at: https://doi.org/10.5281/zenodo.16974644. The artifact contains (1) PickleBall source code, (2) a dataset of malicious and benign models, and (3) survey data of the pickle model ecosystem.
@@ Acknowledgements
We thank our anonymous reviewers for their valuable feedback, and Joseph Lucas, John Irwin, and Suha Hussain for their insightful discussions of pickle security. We thank Beatrice Casey for providing access to the ModelTracer tool (Casey et al., 2024), and to Shenao Wang for sharing malicious models discovered on Hugging Face (Zhao et al., 2024). This work is supported by the National Science Foundation (NSF), through award CNS-2238467. Andreas Kellas is supported by the Department of Defense (DoD) National Defense Science and Engineering (NDSEG) Fellowship program. Yaniv David is supported by the Center for Computer Engineering at the Technion. James Davis is supported by funding from Socket. Junfeng Yang is supported by funding from Google.